OpenStudy (anonymous):

Are y=4x+12 and -4x+3y=21 parallel? Please explain i need an example how to do this.

jigglypuff314 (jigglypuff314):

put -4x + 3y = 21 into slope intercept form: y=mx+b then know that parallel lines have equal slopes so -4x+3y=21 -> y = 4/3 x + 7 (which has a slope of 4/3) y=4x+12 has a slope of 4 4/3 does not equal 4